# Improved polynomial rates of memory loss for nonstationary intermittent dynamical systems

### v1

@@ Abstract.
Date: October 28, 2024
A.K. is supported by EPSRC grant EP/V053493/1. J.L. is supported by JSPS via the project LEADER. The authors thank Wael Bahsoun for useful discussions, as well as Loughborough University and Tokai University for hospitality over the course of this work.
@@ 1. Introduction
[…]
@@
[…]
@@
[…]
@@
${\mathcal{C}}_{*}(\beta)=\{f\in C((0,1])\cap L^{1}\,:\,$ $f\geq 0$, $f$ decreasing,
[…]
@@
$|(T_{1,n})_{*}(\mu-\nu)|=O(n^{-\gamma_{*}}),\quad\text{as $n\to\infty$.}$ (1.3)
[…]
@@
#$\frac{\#\{1\leq k\leq n\>:\>\gamma_{k}\leq\gamma\}}{n}\in[a(1-\kappa),a(1+\kappa)]$ (1.4)
@@ 2.1. Nonstationary nonuniformly expanding dynamical systems, nonuniform tail bounds
We suppose that there exist constants $\lambda>1$, $K>0$, $\delta_{0}>0$ and $n_{0}\geq 1$ such that the following assumptions hold.
@@
    If $a\subset Y_{k}$, then the map $F_{a}=T_{k,k+\tau_{k}(a)-1}:a\to Y_{k+\tau_{k}(a)}$ is a bijection, and for all $y,y^{\prime}\in a$,
@@
    For all $x,x^{\prime}\in a$, with $F_{a}=T_{k,k+\tau_{k}(a)-1}$ as above,
@@
[…]
@@
The proof is similar to in the stationary case [15, Section 4.1]. For completeness, we provide more details in Appendix A. ∎
@@ 3.1. Liverani–Saussol–Vaienti maps
We consider sequences $(T_{n})$ of maps in the family (1.2). First, we prove Theorem 1.1 concerning the rate of memory loss for deterministic compositions, after which we look at applications for random compositions.
@@
$x_{n}(k)=g_{k}\cdots g_{k+n-1}(1)\quad\text{and}\quad y_{n}(k)=hg_{k+1}\cdots g_{k+n-1}(1).$
Then, for all $k\geq 1$ and $n\geq 1$, $T_{k,k+n-1}$ maps both $[x_{n+1}(k),x_{n}(k)]$ and $[y_{n+1}(k),y_{n}(k)]$ bijectively onto $Y$. For each $k\geq 1$, we define a partition of $X$ by
@@
Let $\mu$ be a probability measure with density in ${\mathcal{C}}_{*}(\gamma)$. Then $\mu$ has tail bound $r(n)=O(n^{1-1/\gamma})$ with respect to $T_{1},T_{2},\ldots$ Moreover, for each $k\geq 1$, $h^{k}$ in (2.1) satisfies $h^{k}(\ell)\leq C(1\vee(\ell-\Theta k))^{-1/\gamma}$ where $C$ is a constant independent of $k$, and $\Theta=4\kappa(1-\kappa)^{-1}$.
@@
Let $\gamma_{n_{1}},\gamma_{n_{2}},\ldots$ be the sub-sequence of all parameters $\gamma_{n_{j}}$ with $\gamma_{n_{j}}\leq\gamma$, which by (3.2) must be an infinite sequence. Let $\widetilde{T}_{j}=T_{n_{j}}$. It follows by (3.1) that $\mu$ has tail bound $r(\ell)=O(\ell^{1-1/\gamma})$ with respect to the sequence $\widetilde{T}_{1},\widetilde{T}_{2},\ldots$
@@
$G_{\ell}(k)$ =#$=\#\{k\leq j\leq k+\ell-1\>:\>\gamma_{j}\leq\gamma\}=\#\bigl{(}\{n_{j}\>:\>j\in{\mathbb{N}}\}\cap[k,k+\ell-1]\bigr{)}.$
@@
$\geq\frac{a(1-\kappa)}{2}\biggl{(}\ell-\frac{4\kappa}{1-\kappa}k\biggr{)},$ (3.3)
provided that $\ell\geq 4(1+\kappa)/(1-\kappa)$. ∎
@@
The map preserves an absolutely continuous probability measure if and only if $\gamma\beta<1$. For sequences $T_{1},T_{2},\ldots$ of maps in (3.4) with parameters $\gamma_{k}\leq\gamma_{*}$ and $\gamma_{k}\leq\beta_{*}$ where $\gamma_{*}\beta_{*}<1$, an application of Theorem 2.6 yields the following estimate provided that (3.2) holds: for any two probability measures $\mu$ and $\mu^{\prime}$ with Hölder continuous densities,
@@ 3.1.1. Random compositions
Let $\omega=(\omega_{n})$ be a random sequence of parameters sampled from a probability space $(\Omega,{\mathcal{F}},\mathbf{P})=(\Omega_{0}^{{\mathbb{N}}},{\mathcal{E}}^{{\mathbb{N}}},\mathbf{P})$, where $\Omega_{0}\subset(0,1)$ is an interval, ${\mathcal{E}}$ is the Borel sigma-algebra on $\Omega_{0}$, and =${\mathbb{N}}=\{1,2,\ldots\}$. Denote by $T_{n}\circ\cdots\circ T_{1}$ compositions of maps $T_{k}$ along the random sequence $\omega$, where each $T_{k}$ corresponds to a map with parameter $\omega_{k}$. As a consequence of Theorem 1.1, we obtain the following result:
@@
[…]
@@
    Here, ${\mathcal{F}}_{1}^{i}$ is the sigma-algebra on $\Omega$ generated by the projections $\pi_{1},\ldots,\pi_{i}$; ${\mathcal{F}}_{i+n}^{\infty}$ is the sigma-algebra generated by $\pi_{i+n}\pi_{i+n+1},\ldots$, with $\pi_{k}(\omega)=\omega_{k}$; and
    $\alpha({\mathcal{F}}_{1}^{i},{\mathcal{F}}_{j}^{\infty})=\sup_{A\in{\mathcal{F}}_{1}^{i},\,B\in{\mathcal{F}}_{j}^{\infty}}|\mathbf{P}(A\cap B)-\mathbf{P}(A)\mathbf{P}(B)|.$
@@
Define $\varphi$ as in the proof of Corollary 3.3. By (B3), it is enough to show that, for $\mathbf{P}$-a.e. $\omega\in\Omega$,
@@
[…]
@@
[…]
By (B1), $I\to 0$ as $n\to\infty$. By (B2), $(\xi_{n})$ is a uniformly bounded sequence of $\alpha$-mixing random variables whose $\alpha$-mixing coefficients decay at the rate $O(\log^{-\beta}(n))$ with $\beta>1$. It follows from [26, Corollary 1] that $II\to 0$ as $n\to\infty$, for $\mathbf{P}$-a.e. $\omega\in\Omega$. Thus, we obtain (3.6). ∎
@@ 3.2. Pikovsky maps
For $\gamma>1$, let $T:X\to X$ be a map on $X=[-1,1]$ defined implicitly by the relations
@@
and by letting $T(x)=-T(-x)$ for $x\in[-1,0]$. The graph of $T$ is illustrated in Figure 1. The map has neutral fixed points at $x=1,-1$, while at $x=0$ its derivative becomes infinite. There is an interplay between the neutrality of the fixed points and the degree of the singularity, which is controlled by the parameter $\gamma$:
$T^{\prime}(x)\approx C_{\gamma}x^{1/\gamma-1}-1\quad\text{as $x\searrow 0$},$
@@
Before proving Theorem 3.8, we state a corollary for random compositions. Let $(\omega_{n})$ be a random sequence of parameters sampled from a probability space $(\Omega,{\mathcal{F}},\mathbf{P})=(\Omega_{0}^{{\mathbb{N}}},{\mathcal{E}}^{{\mathbb{N}}},\mathbf{P})$, where $\Omega_{0}=[\gamma_{-},\gamma_{+}]$ with $31<\gamma_{-}<\gamma_{+}<$ 3, and ${\mathcal{E}}$ is the Borel $\sigma$-algebra on $\Omega_{0}$. As in Section 3.1.1, denote by $T_{n}\circ\cdots\circ T_{1}$ compositions of maps $T_{k}$ from (3.7) along the random sequence $\omega$, where each $T_{k}$ corresponds to a map with parameter $\omega_{k}$.
@@
Let $(T_{k})$ be a sequence of maps in the family (3.7) with parameters $\gamma_{k}\in[\gamma_{-},\gamma_{+}]$, where $31<\gamma_{-}<\gamma_{+}<$ 3. Let $X_{-}=(-0,1)$ and $X_{+}=(0,1)$. Let $\lambda$ be the Lebesgue measure on $X$ normalized to probability. Denote by $g_{k,-}:(-1,1)\to X_{-}$ and $g_{k,-}:(-1,1)\to X_{+}$ the left and right inverse branches of $T_{k}$, respectively.
@@
Then $\{\Delta_{n}^{-}(k)\}_{n\geq 1}$ is a (mod $\lambda$) partition of $X_{-}$, $\{\Delta_{n}^{+}(k)\}_{n\geq 1}$ is a (mod $\lambda$) partition of $X_{+}$, and $T_{k}$ maps $\Delta_{n}^{\pm}(k)$ bijectively onto $\Delta_{n-1}^{\pm}(k+1)$. Moreover, $\Delta^{+}_{\ell}(k)=(x_{\ell}^{+}(k),x_{\ell+1}^{+}(k))$ and $\Delta^{-}_{\ell}(k)=(x_{\ell+1}^{-}(k),x_{\ell}^{-}(k))$, where $x_{\ell}^{+}(k)=g_{k,+}\cdots g_{k+\ell-1,+}(0)$.
@@
Then $\tau_{k}$ is constant on the partition elements of ${\mathcal{P}}_{k}$, so that (NU:2) is satisfied. By inspecting the definition of $T$ in (3.7), we see that $|F_{a}(y)-F_{a}(y^{\prime})|\geq\Lambda|y-y^{\prime}|>0$ holds for all $y,y^{\prime}\in a\subset Y_{k}$ whenever $a\in{\mathcal{P}}_{k}$ where $F_{a}=T_{k,k+\tau_{k}(a)-1}$ and $\Lambda=\Lambda(\gamma_{-},\gamma_{+})>1$. Then (NU:3) follows by [20, Lemma 2.11] provided that $31<\gamma_{-}<\gamma_{+}<$ 3:
@@
Assume that $31<\gamma_{-}<\gamma_{+}<$ 3. Then there exists $K>0$ such that for any $k\geq 1$ and $a\subset Y_{k}$, $a\in{\mathcal{P}}_{k}$,
@@
Next, suppose that $z$ and $z^{\prime}$ belong to different smoothness components, say $z\in\delta_{n}^{-}(k)$ and $z^{\prime}\in\delta_{n}^{+}(k)$. In this case we can exploit the trick described in [4, Remark 1]. Namely, by symmetry of $T_{k}$ we have $-z^{\prime}\in\delta_{n}^{-}(k)$ and $T_{j}^{\prime}(y)=T_{j}^{\prime}(-x)$. Since each $T_{j}$ is an odd function,
@@
where the last inequality is due to [4]. Hence, $\lambda(\cup_{\ell\geq n}\Delta_{\ell}^{+}(k))=\lambda((x_{n-1}(k)^{+},1))=O(n^{-1/(\gamma-1)})$. The same estimate holds with $\Delta_{\ell}^{-}(k)$ in place of $\Delta_{\ell}^{+}(k)$, leading to $\lambda(\cup_{\ell\geq n}\Delta_{\ell}(k))=O(n^{-1/(\gamma-1)})$. On the other hand, by [20, Lemma 2.4],
@@
Next, let $(T_{j})$ be a sequence of maps as in Theorem 3.8. Similar to the proof of Theorem 1.1, we let $\widetilde{T}_{j}=T_{n_{j}}$ where $\gamma_{n_{1}},\gamma_{n_{2}},\ldots$ is the subsequence of all parameters $\gamma_{n_{j}}$ with $\gamma_{n_{j}}\leq\gamma$ and, for each $k,\ell\geq 1$, define
$G_{\ell}(k)$ =#$=\#\{k\leq j\leq k+\ell-1\>:\>\gamma_{j}\leq\gamma\}=\#\bigl{(}\{n_{j}\>:\>j\in{\mathbb{N}}\}\cap[k,k+\ell-1]\bigr{)}.$
@@
where $\widetilde{T}_{m(k),m(k)+\ell-1}=\widetilde{T}_{m(k)}\circ\cdots\circ\widetilde{T}_{m(k)+\ell-1}$. By the definitions of $\Delta_{n}(k)$ and $\delta_{n}(k)$, we have
@@
$\rho_{a}(y):=\frac{d(F_{a})_{*}(\mu|_{a})}{dm_{k+\tau_{k}(a)}}(y)\propto\frac{\rho(y_{a})}{F_{a}^{\prime}(y_{a})},$
where $y_{a}=F_{a}^{-1}(y)$. Hence, for $y,z\in Y_{k+\tau_{k}(a)}$, we have
@@
Finally, if $\rho$ is Hölder continuous with exponent $\eta\in(0,1]$, then $\mu$ is regular with respect to the metric $d_{\eta}(x,y)=|x-y|^{\eta}$. Therefore, we can deal with this case by simply changing the metric from $d(x,y)=|x-y|$ to $d_{\eta}$ in the definitions of Section 2.1. ∎

### v2

@@ Abstract.
Date: September 19, 2025
A.K. is supported by EPSRC grant EP/V053493/1. J.L. is supported by JSPS via the project LEADER. The authors thank Wael Bahsoun for useful discussions, Loughborough University and Tokai University for hospitality over the course of this work, and anonymous reviewers for many helpful comments and questions.
@@ 1. Introduction
A classical topic in dynamical systems theory is the decay of correlations, which describes how quickly initial distributions converge toward an invariant distribution. However, a nonstationary dynamical system may lack any natural invariant distribution. In this context, memory loss, a counterpart to correlation decay, was studied by Ott, Stenlund, and Young [22]. Memory is said to be lost if the time-evolutions $\mu_{n},\mu_{n}^{\prime}$ of any two sufficiently regular initial distributions $\mu_{0},\mu_{0}^{\prime}$ on $X$ satisfy $\|\mu_{n}-\mu_{n}^{\prime}\|\to 0$ as $n\to\infty$ with respect to a suitable notion of distance $\|\cdot\|$. This property can be interpreted as all regular measures being attracted by the same moving target in a space of measures. The rate of memory loss has been the subject of several studies in the dynamical systems literature, including [19, 1, 5, 27, 13], and it plays an important role in further statistical analysis of nonstationary models, as evidenced by studies such as [28, 14, 9, 17, 7, 8, 10, 21], among others.
@@
We now briefly describe the contents of our study. Building on the coupling approach as in Korepanov, Kosloff, and Melbourne [15], in a previous study [16] we analyzed the rate of memory loss in nonstationary nonuniformly expanding dynamical systems. We considered an abstract framework where the time-dependent trajectory $T_{1,n}:=T_{n}\circ\cdots\circ T_{1}$ makes frequent returns to a reference set $Y\subset X$, with first return dynamics that exhibit “good” distortion properties. Under a polynomially decaying upper bound on the tails of first return times, we derived memory loss estimates of the form
@@
for suitable probability measures $\mu$ and $\nu$, where $|\cdot|$ is the total variation distance and $(\cdot)_{*}$ is the pushforward. The rate coincides with the sharp rate established by Gouëzel [11] in the stationary case, meaning that it cannot be improved without additional information about the sequence $(T_{n})$. On the other hand, the estimates in [16] represent a worst-case scenario where all maps satisfy uniform tail bounds of the same polynomial order. In contrast, the present manuscript derives estimates on the rate of memory loss assuming non-uniform tail bounds that can vary with time. This flexibility enables us to address questions such as how the rate of convergence in (1.1) is influenced by subsequences $(T_{n_{k}})$ of “good” maps, i.e. maps for which the decay rate of stationary correlations is fast in comparison with other maps in the sequence. For an illustration of this phenomenom, see Theorem 1.1 below.
@@
${\mathcal{C}}_{*}(\beta)=\{f\in C((0,1])\cap L^{1}\,:\,$ $\geq 0, ff$ decreasing,
$\text{$x^{\beta+1}f$ increasing, $f(x)\leq a_{\beta}x^{-\beta}\int_{0}^{1}f(x)\,dx$}\}.$
@@
$|(T_{1,n})_{*}(\mu-\nu)|=O(n^{-1/\gamma_{*}}),\quad\text{as $n\to\infty$.}$ (1.3)
In other words, memory loss occurs at a rate corresponding to the map with parameter $\gamma_{*}$, which is optimal for a general sequence $(T_{n})$ with $\sup_{n}\gamma_{n}\leq\gamma_{*}$. However, the following result shows that if the upper and lower asymptotic densities of “good” maps in the sequence $(T_{1,n})$ are sufficiently close, this rate can be improved:
@@
$\frac{\#\{1\leq k\leq n\>:\>\gamma_{k}\leq\gamma\}}{n}\in[a(1-\kappa),a(1+\kappa)]$ (1.4)
@@
Remark 1.4.
An interesting question, for which we do not have an answer, is whether the assumption on the upper asymptotic density of good maps is necessary in Theorem 1.1. It is important for our arguments, however: it makes sure that the good maps are distributed sufficiently uniformly (see the proof of Proposition 3.1), for which the lower bound alone is not sufficient.
@@ 2.1. Nonstationary nonuniformly expanding dynamical systems, nonuniform tail bounds
We suppose that there exist constants $\lambda>1$ and $K>0$ such that the following assumptions hold.
@@
    If $a\subset Y_{k}$, then the map $F_{a}=T_{k,k+\tau_{k}(a)-1}|_{a}:a\to Y_{k+\tau_{k}(a)}$ is a bijection, and for all $y,y^{\prime}\in a$,
@@
    For all $x,x^{\prime}\in a$, with $F_{a}=T_{k,k+\tau_{k}(a)-1}|_{a}$ as above,
@@
Moreover, for condition (i) it is sufficient that $\sup_{k\geq 1}\int\tau_{k}^{p}\,dm_{k}<\infty$ holds for some $p>1$.
@@
The proof is similar to that in the stationary case [15, Section 4.1]. For completeness, we provide more details in Appendix A. ∎
@@ 3.1. Liverani–Saussol–Vaienti maps
We consider sequences of maps $T_{n}:X\to X$ in the family (1.2) with $X=[0,1]$. First, we prove Theorem 1.1 concerning the rate of memory loss for deterministic compositions, after which we look at applications for random compositions.
@@
$x_{n}(k)=(g_{k}\circ\cdots\circ g_{k+n-1})(1)\quad\text{and}\quad y_{n}(k)=(h\circ g_{k+1}\circ\cdots\circ g_{k+n-1})(1).$
Then, for all $k\geq 1$ and $n\geq 1$, $T_{k,k+n-1}$ maps both $x_{n+1}(k),x_{n}(k)]$ and $y_{n+1}(k),y_{n}(k)]$ bijectively onto $Y$. For each $k\geq 1$, we define a partition of $X$ by
@@
Let $\mu$ be a probability measure with density in ${\mathcal{C}}_{*}(\gamma)$. Then $\mu$ has tail bound $r(n)=O(n^{1-1/\gamma})$ with respect to $T_{1},T_{2},\ldots$ Moreover, for each $k\geq 1$, $h^{k}$ in (2.1) satisfies $h^{k}(\ell)\leq C(1\vee(\ell-\Theta k))^{-1/\gamma}$ where $C$ is a constant independent of $k$, and $\Theta=2\kappa(1-\kappa)^{-1}$.
@@
Let $\gamma_{n_{1}},\gamma_{n_{2}},\ldots$ be the subsequence of all parameters $\gamma_{n}$ with $\gamma_{n}\leq\gamma$, which by (3.2) must be an infinite sequence. Let $\widetilde{T}_{j}=T_{n_{j}}$. It follows by (3.1) that $\mu$ has tail bound $r(\ell)=O(\ell^{1-1/\gamma})$ with respect to the sequence $\widetilde{T}_{1},\widetilde{T}_{2},\ldots$
@@
$G_{\ell}(k)$ $=\#\{k\leq j\leq k+\ell-1\>:\>\gamma_{j}\leq\gamma\}=\#\bigl(\{n_{j}\>:\>j\in{\mathbb{N}}\}\cap[k,k+\ell-1]\bigr).$
@@
$\geq a(1-\kappa)\biggl(\ell-\frac{2\kappa}{1-\kappa}k\biggr).$ (3.3)
@@
The map $T$ preserves an absolutely continuous probability measure if and only if $\gamma\beta<1$. For sequences $T_{1},T_{2},\ldots$ of maps in (3.4) with parameters $\gamma_{k}\leq\gamma_{*}$ and $\gamma_{k}\leq\beta_{*}$ where $\gamma_{*}\beta_{*}<1$, an application of Theorem 2.6 yields the following estimate provided that (3.2) holds: for any two probability measures $\mu$ and $\mu^{\prime}$ with Hölder continuous densities,
@@ 3.1.1. Random compositions
Let $\omega=(\omega_{n})$ be a random sequence of parameters sampled from a probability space $(\Omega,{\mathcal{F}},\mathbf{P})=(\Omega_{0}^{{\mathbb{N}}},{\mathcal{E}}^{{\mathbb{N}}},\mathbf{P})$, where $\Omega_{0}\subset(0,1)$ is an interval, $\mathcal{E}$} is the Borel sigma-algebra on $\Omega_{0}$, and ${\mathbb{N}}=\{1,2,\ldots\}$. Denote by $T_{n}\circ\cdots\circ T_{1}$ compositions of maps $T_{k}$ along the random sequence $\omega$, where each $T_{k}$ corresponds to a map with parameter $\omega_{k}$. As a consequence of Theorem 1.1, we obtain the following result:
@@
Recall that, given two sub-$\sigma$-algebras $\mathcal{U},\mathcal{V}\subset\mathcal{{\mathcal{F}}}$, the $\alpha$-mixing coefficient between $\mathcal{U}$ and $\mathcal{V}$ is defined by
$\alpha(\mathcal{U},\mathcal{V})=\sup_{U\in\mathcal{U},\>V\in\mathcal{V}}|\mathbf{P}(U\cap V)-\mathbf{P}(U)\mathbf{P}(V)|.$
@@
    $\lim_{n\to\infty}\int_{\Omega}n^{-1}\sum_{k=0}^{n-1}g\circ\sigma^{k}\,d\mathbf{P}=\int_{\Omega}g\,d\bar{\mathbf{P}}.$
@@
    Here, ${\mathcal{F}}_{1}^{i}$ is the sigma-algebra on $\Omega$ generated by the projections $\pi_{1},\ldots,\pi_{i}$; ${\mathcal{F}}_{i+n}^{\infty}$ is the sigma-algebra generated by $\pi_{i+n},\pi_{i+n+1},\ldots$, with $\pi_{k}(\omega)=\omega_{k}$.
@@
Define $\varphi=\mathbf{1}_{[0,\gamma]}\circ\pi_{1}$ as in the proof of Corollary 3.3. By (B3), it is enough to show that, for $\mathbf{P}$-a.e. $\omega\in\Omega$,
@@
$I=\int_{\Omega}n^{-1}\sum_{k=0}^{n-1}\varphi\circ\sigma^{k}\,d\mathbf{P}-\int_{\Omega}\varphi\,d\bar{\mathbf{P}}$
@@
$II=n^{-1}\sum_{k=0}^{n-1}\xi_{k},\quad\xi_{k}=\varphi\circ\sigma^{k}-\int_{\Omega}\varphi\circ\sigma^{k}\,d\mathbf{P}.$
By (B1), $I\to 0$ as $n\to\infty$. Moreover, $\sup_{k}\sup_{\omega\in\Omega}|\xi_{k}(\omega)|\leq 2$ and, by (B2),
$\alpha(\sigma(\xi_{i}:0\leq i\leq j),\sigma(\xi_{i}:i\geq j+n))\leq\alpha(n)=O(\log^{-\beta}(n)).$
Here, $\sigma(\xi_{i}:0\leq i\leq j)$ and $\sigma(\xi_{i}:i\geq j+n)$ denote the sub-$\sigma$-algebras generated by $\{\xi_{i}\}_{0\leq i\leq j}$ and $\{\xi_{i}\}_{i\geq j+n}$, respectively. Thus, $(\xi_{n})$ is a uniformly bounded sequence of $\alpha$-mixing random variables with coefficients decaying at the rate $O(\log^{-\beta}(n))$ with $\beta>1$. It follows from [26, Corollary 1] that $II\to 0$ as $n\to\infty$, for $\mathbf{P}$-a.e. $\omega\in\Omega$. Thus, we obtain (3.6). ∎
@@ 3.2. Pikovsky maps
For $\gamma>1$, let $T:X\to X$ be a map on X=[1,1]X=[-1,1] defined implicitly by the relations
@@
and by letting T(x)=T(x)T(x)=-T(-x) for $x\in[-1,0]$. The graph of $T$ is illustrated in Figure 1. The map has neutral fixed points at x=1,1x=1,-1, while at $x=0$ its derivative becomes infinite. There is an interplay between the neutrality of the fixed points and the degree of the singularity, which is controlled by the parameter $\gamma$:
$T^{\prime}(x)\approx C_{\gamma}x^{1/\gamma-1}\quad\text{as $x\searrow 0$},$
@@
Before proving Theorem 3.8, we state a corollary for random compositions. Let $(\omega_{n})$ be a random sequence of parameters sampled from a probability space $(\Omega,{\mathcal{F}},\mathbf{P})=(\Omega_{0}^{{\mathbb{N}}},{\mathcal{E}}^{{\mathbb{N}}},\mathbf{P})$, where $\Omega_{0}=[\gamma_{-},\gamma_{+}]$ with $1<\gamma_{-}<\gamma_{+}<3$, and $\mathcal{E}$} is the Borel $\sigma$-algebra on $\Omega_{0}$. As in Section 3.1.1, denote by $T_{n}\circ\cdots\circ T_{1}$ compositions of maps $T_{k}$ from (3.7) along the random sequence $\omega$, where each $T_{k}$ corresponds to a map with parameter $\omega_{k}$.
@@
Let $(T_{k})$ be a sequence of maps in the family (3.7) with parameters $\gamma_{k}\in[\gamma_{-},\gamma_{+}]$, where $1<\gamma_{-}<\gamma_{+}<3$. Let $X_{-}=(-0,1)$ and $X_{+}=(0,1)$. Let $\lambda$ be the Lebesgue measure on $X$ normalized to probability. Denote by $g_{k,-}:(-1,1)\to X_{-}$ and $g_{k,-}:(-1,1)\to X_{+}$ the left and right inverse branches of $T_{k}$, respectively.
@@
Then $\{\Delta_{n}^{-}(k)\}_{n\geq 1}$ is a (mod $\lambda$) partition of $X_{-}$, $\{\Delta_{n}^{+}(k)\}_{n\geq 1}$ is a (mod $\lambda$) partition of $X_{+}$, and $T_{k}$ maps $\Delta_{n}^{\pm}(k)$ bijectively onto $\Delta_{n-1}^{\pm}(k+1)$. Moreover, $\Delta^{+}_{\ell}(k)=(x_{\ell}^{+}(k),x_{\ell+1}^{+}(k))$ and $\Delta^{-}_{\ell}(k)=(x_{\ell+1}^{-}(k),x_{\ell}^{-}(k))$, where $x_{\ell}^{+}(k)=(g_{k,+}\circ\cdots\circ g_{k+\ell-1,+})(0)$ and $x_{\ell}^{-}(k)=(g_{k,-}\circ\cdots\circ g_{k+\ell-1,-})(0)$.
@@
Then $\tau_{k}$ is constant on the partition elements of ${\mathcal{P}}_{k}$, so that (NU:2) is satisfied. By inspecting the definition of $T$ in (3.7), we see that $|F_{a}(y)-F_{a}(y^{\prime})|\geq\Lambda|y-y^{\prime}|>0$ holds for all $y,y^{\prime}\in a\subset Y_{k}$ whenever $a\in{\mathcal{P}}_{k}$ where $F_{a}=T_{k,k+\tau_{k}(a)-1}$ and $\Lambda=\Lambda(\gamma_{-},\gamma_{+})>1$. Then (NU:3) follows by [20, Lemma 2.11] provided that $1<\gamma_{-}<\gamma_{+}<3$:
@@
Assume that $1<\gamma_{-}<\gamma_{+}<3$. Then there exists $K>0$ such that for any $k\geq 1$ and $a\subset Y_{k}$, $a\in{\mathcal{P}}_{k}$,
@@
Next, suppose that $z$ and $z^{\prime}$ belong to different smoothness components, say $z\in\delta_{n}^{-}(k)$ and $z^{\prime}\in\delta_{n}^{+}(k)$. In this case we can exploit the trick described in [4, Remark 1]. Namely, by symmetry of $T_{k}$ we have $z^{\prime}\in\delta_{n}^{-}(k)$ and $T_{j}^{\prime}(y)=T_{j}^{\prime}(-x)$. Since each $T_{j}$ is an odd function,
@@
where the last inequality is due to [4]. Hence, $\lambda(\cup_{\ell\geq n}\Delta_{\ell}^{+}(k))=\lambda((x_{n-1}^{+}(k),1))=O(n^{-1/(\gamma-1)})$. The same estimate holds with $\Delta_{\ell}^{-}(k)$ in place of $\Delta_{\ell}^{+}(k)$, leading to $\lambda(\cup_{\ell\geq n}\Delta_{\ell}(k))=O(n^{-1/(\gamma-1)})$. On the other hand, by [20, Lemma 2.4],
@@
Next, let $(T_{j})$ be a sequence of maps as in Theorem 3.8. Similar to the proof of Theorem 1.1, we let $\widetilde{T}_{j}=T_{n_{j}}$ where $\gamma_{n_{1}},\gamma_{n_{2}},\ldots$ is the subsequence of all parameters $\gamma_{n}$ with $\gamma_{n}\leq\gamma$ and, for each $k,\ell\geq 1$, define
$G_{\ell}(k)$ $=\#\{k\leq j\leq k+\ell-1\>:\>\gamma_{j}\leq\gamma\}=\#\bigl(\{n_{j}\>:\>j\in{\mathbb{N}}\}\cap[k,k+\ell-1]\bigr).$
@@
where $\widetilde{T}_{m(k),m(k)+\ell-1}=\widetilde{T}_{m(k)+\ell-1}\circ\cdots\circ\widetilde{T}_{m(k)}$. By the definitions of $\Delta_{n}(k)$ and $\delta_{n}(k)$, we have
@@
$\rho_{a}(y):=\frac{d(F_{a})_{*}(\mu|_{a})}{dm_{k+\tau_{k}(a)}}(y)=C_{k,a}\frac{\rho(y_{a})}{F_{a}^{\prime}(y_{a})},$
where $y_{a}=F_{a}^{-1}(y)$ and $C_{k,a}$ is a scaling constant. Hence, for $y,z\in Y_{k+\tau_{k}(a)}$, we have
@@
Finally, if $\rho$ is Hölder continuous with exponent $\eta\in(0,1]$, then $\mu$ is regular with respect to the metric $d_{\eta}(x,y)=|x-y|^{\eta}$. Therefore, we can deal with this case by simply changing the metric from d(x,y)=|xy|d(x,y)=|x-y| to $d_{\eta}$ in the definitions of Section 2.1. ∎